# CCIS-Diff: A Generative Model with Stable Diffusion Prior for Controlled Colonoscopy Image Synthesis

### v1

@@ 1 Introduction
Colonoscopy is an essential tool for detecting adenomatous polyps and reducing rectal cancer mortality rates [1, luo2024codeswap, ma2024generative]. However, training models for automatic polyp detection is challenging due to the small scale of available colonoscopy datasets, making it difficult to have sufficient robustness and generalization that meet real-world clinical demands.
@@ 2.1 Multi-Modal Colonoscopy Dataset
A key challenge for current diffusion models in colonoscopy image synthesis [2, 6] is the absence of a dataset with consistent spatial characteristics of polyps and their corresponding clinical textual descriptions. Existing datasets, such as EndoScene [9], CVC-ClinicDB [10], and Kvasir [11], primarily contain image-mask pairs that focus on polyp regions. However, these datasets do not include detailed clinical text descriptions, limiting their usefulness for training models that aim to generate clinically accurate and diverse synthetic images. To address this limitation, we construct a novel multi-modal colonoscopy dataset that integrates three essential components: colonoscopy images, segmentation masks, and clinical text descriptions. Such a dataset is vital for fine-tuning pretrained diffusion models, allowing them to adapt effectively to controlled colonoscopy image synthesis.
@@ 3.2 Baseline Algorithms and Evaluation Metrics
Table 3: Comparisons of different settings applied on two polyp segmentation baselines [1, 15] across three public datasets [9, 10, 11].

### v2

@@ 1 Introduction
Colonoscopy is an essential tool for detecting adenomatous polyps and reducing rectal cancer mortality rates [1]. However, training models for automatic polyp detection is challenging due to the small scale of available colonoscopy datasets, making it difficult to have sufficient robustness and generalization that meet real-world clinical demands.
@@ 2.1 Multi-Modal Colonoscopy Dataset
A key challenge for current diffusion models in colonoscopy image synthesis [2, 6] is the absence of a dataset with consistent spatial characteristics of polyps and their corresponding clinical textual descriptions. Existing datasets, such as EndoScene [9], CVC-ClinicDB [10], and Kvasir [11], primarily contain image-mask pairs that focus on polyp regions. However, these datasets do not include detailed clinical text descriptions, limiting their usefulness for training models that aim to generate clinically accurate and diverse synthetic images. To address this limitation, we built a new multi-modal colonoscopy dataset based on existing datasets, containing three key components: colonoscopy images, segmentation masks, and clinical text descriptions. Such a dataset is vital for fine-tuning pretrained diffusion models, allowing them to adapt effectively to controlled colonoscopy image synthesis.
@@ 3.2 Baseline Algorithms and Evaluation Metrics
Table 3: Comparisons of different settings applied on two polyp segmentation baselines across three public datasets.